# ZS-VCOS: Zero-Shot Outperforms Supervised Video Camouflaged Object Segmentation

### v1

@@ 4.1.1 Metrics and Datasets
In this paper, we examine two variants of the MoCA dataset [21]: MoCA-Mask [9] and MoCA-Filtered [56]. MoCA-Mask is a segmentation dataset, and our evaluation approach aligns with SLT-Net [9]. We report the following metrics: S-measure ($S_{a}$) [8], weighted F-measure ($F_{\beta}^{w}$) [28], and Mean Absolute Error (MAE). More details on these metrics can be found in the SLT-Net paper and their original sources. We did not use E-measure [11], mean Dice coefficient, and mean Intersection-over-Union (IoU) in our comparison with previous methods because they depend on metric calculation methods, which we will explain later. For our internal comparisons in the ablation study, we primarily focus on S-measure, E-measure, and mean IoU.
Metrics such as mean IoU can be computed in two primary ways: (1) calculating IoU for each frame individually, averaging across frames, and then averaging across videos, or (2) continuously updating a confusion matrix across all frames and calculating a final IoU based on cumulative results. The second method addresses scenarios where predictions and ground truths might both be empty for certain frames. However, in the MoCA-Mask dataset, all frames contain positive pixels, making this issue irrelevant. Despite this, the two calculation methods can yield different results, occasionally significantly so. To ensure consistency and fairness, we adopted the evaluation code and methodology from SLT-Net, which calculates metrics by first averaging across frames and then averaging across videos. This methodology might differ from other reported methods. We re-calculated these metrics using the SLT-Net evaluation implementation for the supervised state-of-the-art method, ZoomNeXt, and it resulted in slightly elevating its $E_{m}$, mDice, and mIoU compared to original reporting. However $S_{a}$, $F_{\beta}^{w}$, and MAE remains unchanged. Thus, in this paper, we focus on these 3 metrics when comparing them across previous methods.
Although there are forum discussions about this issue [17], there is a lack of formal publications about this issue ([49] mentioned and investigated this problem). We encourage future research to clearly specify their metric calculation methodology and consider adopting this standardized frame-then-video averaging approach to facilitate fair comparisons.
Since the SLT-Net method produces soft outputs with continuous pixel values, they considered multiple thresholds. However, as our method produces binary outputs, we used a single threshold value of 0.5.
@@ 4.1.3 Settings
Since our method is zero-shot without training, we directly evaluated it on the testing set. For MoCA-Mask, to minimize overfitting hyperparameters on the limited test set while ensuring reasonable performance, we slightly adjusted hyperparameters manually and swept the VLM threshold (following [15]) from 0.03 to 0.13 (inclusive) at a step of 0.02. For MoCA-Filtered, we employed a fixed threshold of 0.12 tuned by hand. Optical flow was computed using RAFT-Things [43], employing the implementation provided by Motion Grouping [56] with only forward flow and a frame gap of 1. All input images were passed directly to the model processor without resizing or cropping. The momentum parameter ($m$) was set to 0.9. We used Owlv2-Base-Patch16-Ensemble and Sam2.1-Hiera-Small as our VLM and segmentation models.
@@
Table 3: Success rate (SR) of detection success rate on MoCA-Filtered [56] “SV” stands for supervised training, “SV” stands for self-supervised, and “ZS” stands for zero-shot. Although the previous three methods are trained on VOS datasets, they are not trained on camouflage object datasets. Our method is not trained on any VOS or camouflage datasets and resulted in the highest SR.
@@ 4.2.2 MoCA-Filtered
Our results for MoCA-Filtered are presented in Table 3. Our method outperforms Flow-SAM [52] and Motion Grouping [56], which are trained on non-camouflage video segmentation datasets using supervised and unsupervised approaches, respectively. We achieve a detection success rate of 0.697, compared to 0.645 for Flow-SAM and 0.484 for Motion Grouping.
@@ 4.3 Video-Level Results
We examined individual results for each video in the test set. Quantitative results for mIoU are presented in Table 6, and qualitative results are shown in Figure 3. Both results indicate that our method succeeded in stick_insect_1 and snow_leopard_10, where ZoomNeXt completely failed. Additionally, our approach successfully captured the target in ibex, whereas the other two methods missed it. In arctic_fox_3, our method achieved a significantly higher $F_{\beta}^{w}$ score and effectively avoided stationary objects. Although our method struggled in pygmy_seahorse_0, neither ZoomNeXt nor SLT-Net performed well in this case. In other cases where other methods outperformed ours, the margin of improvement was minimal.
(a) 0.570 0.623 0.205
(b) 0.749 0.868 0.501
(c) 0.776 0.878 0.550
(d) 0.776 0.878 0.550
Table 4: Effects of different VLM prompting strategies
SAM-2 Prompt $S_{\alpha}$ $E_{\phi}$ mIoU
(a) Box Only 0.776 0.873 0.540
(b) Point + Box 0.776 0.878 0.550
Table 5: Effects of different SAM-2 prompts
@@ 4.4 Ablation Study
For our ablation study, we designed 9 configurations, as shown in Table 2. Configuration (a) is a minimum baseline with only object detection, used to compare with prior methods such as LLaVA/Shikra + SAM2 [58]. Our result ($S_{\alpha}=0.621$) is nearly identical to theirs ($S_{\alpha}=0.622$). In (b), we add bidirectional tracking to (a), which slightly improves $S_{\alpha}$ by +0.022 and mIoU by +0.049. In (c), we add optical flow and mean subtraction on top of (b), leading to a significant improvement: $S_{\alpha}$ increases by +0.109 and mIoU by +0.207. In (d), we introduce momentum to (c), resulting in a very small drop in $S_{\alpha}$ (-0.002) but a minor gain in mIoU (+0.005). In (e), we use both optical flow and background subtraction based on camera movements, along with mean subtraction for optical flow. Resulted in a slight increase compared to (d), +0.009 in $S_{\alpha}$ and 0.009 in mIoU. In (f) and (g), we remove tracking entirely to assess its impact. Both show a substantial performance drop, especially in mIoU, indicating that tracking is essential. (g) includes momentum, while (f) does not. In (h), we test forward-only tracking instead of bidirectional. It performs worse than Ours (-0.029 in $S_{\alpha}$ and -0.053 in mIoU), showing bidirectional tracking is more effective. Finally, we removed mean subtraction from Ours, as shown in (i), this resulted in s small increase in $S_{\alpha}$ and mIoU but a lower $E_{\phi}$. This means that subtraction has a minimum impact on performance. When compared to other methods, our method without mean subtraction [(i)] achieved the highest $S_{\alpha}$ and mIoU, while ours achieved the highest $E_{\phi}$ with $S_{\alpha}$ and mIoU very close to (i).
@@
Table 6: mIoU of Each Video in The MoCA-Mask Test Set
@@ 4.5.1 Prior Knowledge in Prompt
In our methods, we explicitly instructed the vision-language model to look for animals or insects highlighted in blue. However, to evaluate the model’s generalization capability, we replaced the specific terms “animal or insect” in the prompts with the more generic term ”object.” The corresponding results are shown in Table 4 row (a). As observed, performance drops significantly when switching from specific categories (animal/insect) to a general object category. We suspect this is because multiple objects are often moving within the video, making it unclear to the model which object it should focus on.
@@ 4.5.2 Mentioning Highlight Color in Prompt
To emphasize motion within the frame, we highlighted the relevant areas in blue. Our ablation study demonstrated the effectiveness of the highlighting itself. Additionally, we explicitly guided the visual language model (VLM) by prompting it to detect animals or insects highlighted in blue. The impact of this prompt was tested by removing it, as shown in row (c) of Table LABEL:tab:_prompt, where the prompt was simply ”an animal or insect.” Without explicitly instructing the model to focus on highlighted areas (indicative of motion), we observed a slight performance drop across all metrics. Nonetheless, performance remained relatively high, suggesting that highlighting motion regions alone aids detection, even without explicit prompting (see comparison with row (b), no highlighting and no instruction for highlighting, in Table 2).
@@ 4.5.3 Negative Prompt
We hypothesized that negative prompts could help the model avoid negative objects. However, as shown in Table 4 row (c), the results are identical to the setting with negative prompts. (Noted that although these two settings achieved the same best performances, their results are not identical at all VLM threshold settings.) This shows that the VLM used (Owlv2 [31]) can effectively avoid non-targeted interest without negative prompting.
4.6 Data Contamination Concerns
When using large foundational models, data contamination is a valid concern. However, OWLv2 is trained using pseudo-labels from WebLI [6] generated by OWL-ViT [30]. This is less concerning, as OWL is primarily trained on image-text pairs without localization information. The detection data used in OWL-Vit was sourced from Object365 [39], Open Images [2], and Visual Genome [18], all of which were published before the original MoCA [21] dataset. Therefore, while contamination is a consideration, it is limited. At the time of this paper’s publication, two new datasets [45, 14] (not yet released) are expected to provide uncontaminated data, and we encourage future work to evaluate our method on these datasets.
5 Conclusion
We introduced ZS-VCOS, a zero-shot method for video camouflaged object segmentation, integrating optical flow, vision-language models, and SAM. Our approach significantly outperformed existing methods, increasing mIoU on the MoCA-Mask dataset from 0.273 to 0.561 and improving detection success on MoCA-Filtered from 0.628 to 0.697. Our findings highlight the potential of zero-shot pipelines for effectively handling camouflaged objects, particularly beneficial in scenarios lacking labeled data.
Acknowledgement
This work was supported by NFRF GR024473 and CFI GR024801. We also thank Weathon Software (https://weasoft.com) for providing computing credits via Google Colab.

### v2

@@ 4.1.1 Metrics and Datasets
In this paper, we examine two variants of the MoCA dataset [22]: MoCA-Mask [9] and MoCA-Filtered [57]. The Camouflage Animal Dataset (CAD) [3] is not used due to its inaccessibility, as the server is offline and the dataset is not provided by a third party. MoCA-Mask is a segmentation dataset, and our evaluation approach aligns with SLT-Net [9]. We report the following metrics: S-measure ($S_{a}$) [8], weighted F-measure ($F_{\beta}^{w}$) [29], and Mean Absolute Error (MAE). More details on these metrics can be found in the SLT-Net paper and their original sources. We did not focus on E-measure [11], mean Dice coefficient, and mean Intersection-over-Union (IoU) in our comparison with previous methods because they will yield different results depending on metric calculation methods, which we will explain in the supplementary material Section LABEL:sec:metric. All standard metrics are provided in Table 1, with metrics that could be miscalculated by the previous method colored in gray. For our internal comparisons in the ablation study, we primarily focus on a subset of these metrics (using a subset of internal comparisons is used in SLT-Net): S-measure, E-measure, and mean IoU. To ensure consistency and fairness, we adopted the evaluation code and methodology from SLT-Net.
Although there are forum discussions about this issue [18], there are only a few publications mentioned about this issue [50, 16]. We encourage future research to clearly specify their metric calculation methodology and consider adopting this standardized frame-then-video averaging approach to facilitate fair comparisons.
Since the SLT-Net method produces soft outputs with continuous pixel values, they considered multiple thresholds and report a max and mean metric. However, as our method produces binary outputs, we used a single threshold value of 0.5.
@@ 4.1.3 Settings
Since our method is zero-shot without training, we directly evaluated it on the testing set. For MoCA-Mask, to minimize overfitting hyperparameters on the limited test set while ensuring reasonable performance in the real world, we slightly adjusted hyperparameters manually and swept the VLM threshold (following [16]) from 0.03 to 0.13 (inclusive) at a step of 0.02. For MoCA-Filtered, we employed a fixed threshold of 0.12 tuned by hand. Optical flow was computed using RAFT-Things [44], employing the implementation provided by Motion Grouping [57] with only forward flow and a frame gap of 1. All input images were passed directly to the model processor without resizing or cropping. The momentum parameter ($m$) was set to 0.9. We used Owlv2-Base-Patch16-Ensemble and Sam2.1-Hiera-Small as our VLM and segmentation models.
@@
Table 3: Success rate (SR) of detection success rate on MoCA-Filtered [57]. “SV” stands for supervised training, “SV” stands for self-supervised, and “ZS” stands for zero-shot. Although the previous three methods are trained on VOS datasets, they are not trained on camouflage object datasets. Our method is not trained on any VOS or camouflage datasets and resulted in the highest SR.
@@ 4.2.2 MoCA-Filtered
Our results for MoCA-Filtered are presented in Table 3. Our method outperforms Flow-SAM [53] and Motion Grouping [57], which are trained on non-camouflage video segmentation datasets using supervised and unsupervised approaches, respectively. We achieve a detection success rate of 0.697, compared to 0.645 for Flow-SAM and 0.484 for Motion Grouping. The improvement here is not as significant as in MoCA-Mask, which may indicate that combining SAM and optical flow, as done in FlowSAM, is already an effective approach.
@@ 4.3 Video-Level Results
We examined individual results for each video in the test set. Quantitative results for mIoU are presented in Table 4 in supplemental material, and qualitative results are shown in Figure 3. Both results indicate that our method succeeded in stick_insect_1 and snow_leopard_10, where ZoomNeXt completely failed. Additionally, our approach successfully captured the target in ibex, whereas the other two methods missed it. In arctic_fox_3, our method achieved a significantly higher $F_{\beta}^{w}$ score and effectively avoided stationary objects. Although our method struggled in pygmy_seahorse_0, neither ZoomNeXt nor SLT-Net performed well in this case. In other cases where other methods outperformed ours, the margin of improvement was minimal.
@@ 4.4 Ablation Study
For our ablation study, we designed 9 configurations, as shown in Table 2. Configuration (a) is a minimum baseline with only object detection, used to compare with prior methods such as LLaVA/Shikra + SAM2 [59]. Our result ($S_{\alpha}=0.621$) is nearly identical to theirs ($S_{\alpha}=0.622$). In (b), we add bidirectional tracking to (a), which slightly improves $S_{\alpha}$ by +0.022 and mIoU by +0.049. In (c), we add optical flow and mean subtraction on top of (b), leading to a significant improvement: $S_{\alpha}$ increases by +0.109 and mIoU by +0.207. In (d), we introduce momentum to (c), resulting in a very small drop in $S_{\alpha}$ (-0.002) but a minor gain in mIoU (+0.005). In (e), we use both optical flow and background subtraction based on camera movements, along with mean subtraction for optical flow, resulted in a slight increase compared to (d), +0.009 in $S_{\alpha}$ and 0.009 in mIoU. In (f) and (g), we remove tracking entirely to assess its impact. Both show a substantial performance drop, especially in mIoU, indicating that tracking is essential. (g) includes momentum, while (f) does not. In (h), we test forward-only tracking instead of bidirectional. It performs worse than Ours (-0.029 in $S_{\alpha}$ and -0.053 in mIoU), showing bidirectional tracking is more effective. Finally, we remove mean subtraction from Ours, as shown in (i), which resulted in a slight increase in $S_{\alpha}$ and mIoU but a lower $E_{\phi}$. This means that subtraction has a minimum impact on performance. This might be due to the limited number of videos in the dataset featuring camera motion with relatively static objects, or because the pipeline relies more effectively on contrast rather than absolute color for object identification. Compared to other methods, our approach without mean subtraction (i) achieved the highest $S_{\alpha}$ and mIoU scores. However, our full method obtained the highest $E_{\phi}$, with $S_{\alpha}$ and mIoU scores close to those of (i).
@@
4.5 Prompting Experements
In supplementary material Section 6.1, we studied the effects of different Owlv2 prompts and SAM-2 prompts. Results show that when given prior knowledge and the color of the highlight to Owlv2, the detection performs the best, and when given boxes and points as prompts to SAM-2, the segmentation performs the best. In that section, we also argued why using prior knowledge is a fair comparison with previous methods.
4.6 Data Contamination Concerns
When using large foundational models, data contamination is a valid concern. However, we examined the training data timeline of Owlv2 and concluded that the contamination from the MoCA dataset is highly limited. A detailed explanation can be found in the supplementary material Section 5.1.
5 Conclusion
We introduced ZS-VCOS, a zero-shot method for video camouflaged object segmentation, integrating optical flow, vision-language models, and SAM. Our approach significantly outperformed existing methods, increasing mIoU on the MoCA-Mask dataset from 0.273 to 0.561 and improving detection success on MoCA-Filtered from 0.628 to 0.697. Our findings highlight the potential of zero-shot pipelines for effectively handling camouflaged objects, particularly beneficial in scenarios lacking labelled data. Our modular design enables easy replacement of improved modules at any pipeline stage, enhancing overall performance.
Our method has several limitations. First, it is designed for videos containing one and only one object. In multi-object scenarios, the tracking and matching components would require modification to handle multiple object associations. Second, the approach relies on a textual description of the target object. While this is significantly less costly than collecting annotated training data, generating an accurate and unambiguous prompt can still be non-trivial in some cases. Potential solutions include incorporating few-shot object detection using example image embeddings from VLM as queries, or integrating an image-to-text captioning tool to automatically generate prompts from reference frames.
Acknowledgement
This work was supported by NFRF GR024801 and CFI GR024473. We also thank Weathon Software (https://weasoft.com) for providing computing credits via Google Colab.
Supplementary Material
5.1 Data Contamination Concerns
When using large foundational models, data contamination is a valid concern. However, OWLv2 is trained using pseudo-labels from WebLI [6] generated by OWL-ViT [31]. This is less concerning, as OWL is primarily trained on image-text pairs without localization information. The detection data used in OWL-Vit was sourced from Object365 [40], Open Images [2], and Visual Genome [19], all of which were published before the original MoCA [22] dataset. Therefore, while contamination is a consideration, it is highly limited. At the time of this paper’s publication, two new datasets [46, 14] (not yet released) could provide uncontaminated data, and we encourage future work to evaluate our method on these datasets.
Data contamination is an acknowledged issue when using large-scale foundational models. However, in our case, while it is possible that visual content from MoCA may have appeared in pretraining corpora, it is highly unlikely that the segmentation ground truth or specific frame-level annotations were included. Therefore, even under worst-case assumptions, the problem reduces to a transductive inference setting. Our pipeline remains zero-shot in the sense that no ground truth of the target task is used in any training stage.
6 Previous Metrics Inconsistance
Metrics such as mean IoU can be computed in three primary ways: (1) calculating IoU for each frame individually, averaging across frames within one video, and then averaging across videos, (2) calculating IoU for each frame, and averaging all frames’ results, or (3) calculating an IoU for all frames, which is equivalent to treating the entire video sequence as a single, large concatenated image for both the predicted masks and the ground truth masks, and then computing the IoU on these two large, combined frames. The three calculation methods can yield different results, occasionally significant. Additionally, the calculation script for each paper varies slightly, such as SLT-Net omits the last frame to keep for flow-based methods. In our paper, we computed the metrics using the SLT-Net script. This methodology might differ from other reported methods. We recalculated these metrics using the SLT-Net evaluation implementation for the supervised state-of-the-art method, ZoomNeXt, and it resulted in slightly elevated results of $E_{m}$=0.755, mDice=0.511, and mIoU=0.438 compared to the original reporting (see Table 1). However $S_{a}$, $F_{\beta}^{w}$, and MAE remains unchanged. Thus, in this paper, we focus on these 3 metrics when comparing them across previous methods.
@@ 6 Previous Metrics Inconsistance
Table 4: mIoU of each video in the MoCA-Mask test set.
(a) 0.570 0.623 0.205
(b) 0.749 0.868 0.501
(c) 0.776 0.878 0.550
(d) 0.776 0.878 0.550
Table 5: Effects of different VLM prompting strategies
SAM-2 Prompt $S_{\alpha}$ $E_{\phi}$ mIoU
(a) Box Only 0.776 0.873 0.540
(b) Point + Box 0.776 0.878 0.550
Table 6: Effects of different SAM-2 prompts
@@ 6.1.1 Prior Knowledge in Prompt
In our methods, we explicitly instructed the vision-language model to look for animals or insects highlighted in blue. However, to evaluate the model’s generalization capability, we replaced the specific terms “animal or insect” in the prompts with the more generic term ”object.” The corresponding results are shown in Table 5 row (a). As observed, performance drops significantly when switching from specific categories (animal/insect) to a general object category. We suspect this is because multiple objects are often moving within the video, making it unclear to the model which objects it should focus on.
We do not consider the use of prompts mentioning animals to disqualify our method as zero-shot. Camouflaged videos often contain multiple moving or camouflaged elements—such as leaves, lighting, or branches—making it ambiguous for a model to determine which object should be segmented without explicit guidance. The model cannot “mind-read” our purpose of the current test. For example, if we are now looking for non-animal objects in the image, the model has no way of knowing this information. Providing a very general and high-level prior (e.g., “animal or insect”) is essential for disambiguating the target in the absence of supervision. Previous work claiming zero-shot, like [59], also used a similar prior in their prompt (“Please provide the coordinates of the bounding box where the animal is camouflaged in the picture”). Previous work that has been trained on MoCA-Mask could effectively learn this information from the dataset, and previous work that has not been trained on MoCA-Mask (Like FlowSAM [53] and Motion Grouping [57]) has been trained to identify the center, large moving object. These methods introduced the prior knowledge by training, making it fair to compare against our method with prompt prior knowledge. Moreover, these models with fixed prior knowledge might be harder to transfer to other domains (e.g., non-animal objects or videos without a center and big objects) without finetuning. We chose the animal dataset MoCA because it is currently the only large-scale, publicly available dataset for video camouflage segmentation. Other datasets that include non-animal camouflaged objects, such as MSVCOD [15], have not been released at the time of this work. Future work testing the generalizability of these methods and our methods is needed.
@@ 6.1.2 Mentioning Highlight Color in Prompt
To emphasize motion within the frame, we highlighted the relevant areas in blue. Our ablation study demonstrated the effectiveness of the highlighting itself. Additionally, we explicitly guided the visual language model (VLM) by prompting it to detect animals or insects highlighted in blue. The impact of this prompt was tested by removing it, as shown in row (c) of Table 5, where the prompt was simply ”an animal or insect.” Without explicitly instructing the model to focus on highlighted areas (indicative of motion), we observed a slight performance drop across all metrics. Nonetheless, performance remained relatively high, suggesting that highlighting motion regions alone aids detection, even without explicit prompting (see comparison with row (b), no highlighting and no instruction for highlighting, in Table 2).
@@ 6.1.3 Negative Prompt
We hypothesized that negative prompts could help the model avoid negative objects. However, as shown in Table 5 row (c), the results are identical to the setting with negative prompts. (Note that although these two settings achieved the same best performances, their results are not identical at all VLM threshold settings.) This shows that the VLM used (Owlv2 [32]) can effectively avoid non-targeted interest without negative prompting.